# On the approximability of random-hypergraph MAX-3-XORSAT problems with quantum algorithms

### v1

@@ III Spectrally folded quantum optimization algorithms
A good choice of $A$ is important for spectrally folded quantum optimization to succeed; if $A$ is chosen to be too close to 1, then we risk failing to well-approximate $H_{P}$ due to the interference effects mentioned in section II.3. A choice of $A$ which is too small will return a suboptimal approximation ratio, and if $A$ is too close to zero, cause instabilities from having a poorly defined problem to solve. Fortunately, for random hypergraph MAX-3-XORSAT instances–and here random refers to the graph itself and not, critically, on how many constraints are satisfied in the ground state–we can predict the threshold $A$ for which we expect a polynomial depth circuit to return states with $E\simeq AE_{\mathrm{GS}}$ from first principles. The ideal value of $A$ depends both on the problem class and on the variation of spectral folding employed; for MAX-3-XORSAT, $A\geq 0.6$ is achievable as derived below in section IV.4. This is a significant leap over the best known classical approximation algorithm for this problem Barak et al. (2015), which offers a weaker guarantee with much more restricted viability, in the worst case, equivalent to $A\to 0$ in our notation. It is also a significant leap over recent quantum approaches to this problem Anshu et al. (2021); Marwaha and Hadfield (2022). We note that minimizing $\left(H-E\right)^{2}$ is not itself a novel idea and has been used in classical and quantum algorithms for finding states close to specific energies in chemical and many-body systems Wang and Zunger (1994a, b); McClean et al. (2016); Santagati et al. (2018); Zhang et al. (2021); Tilly et al. (2022); Tazi and Thom (2023). To our knowledge, however, the use of spectral folding for approximate optimization of CSPs is novel, both in concept and in the analysis we present below to choose $A$ and understand at a deeper level why it presents significant advantages over optimizing the problem $H_{P}$ directly.
@@
From hereon, we let $H_{\mathrm{fold}}$ be the spectrally folded problem Hamiltonian. Having defined it, there are a number of ways we can attempt to find its ground states. The simplest choice, and one that performs well, is AQC/QAOA-inspired state preparation, where we interpolate in Trotterized evolution between the transverse field driver and the folded problem over a time $t_{f}$:
@@
This prescription, with the quadratic folding choice in Eq. (5), is the most straightforward to benchmark using standard quantum simulation packages as each call requires $O\left(N_{C}^{2}\right)$ multiqubit $Z_{i}Z_{j}Z_{k}Z_{l}...$ rotations. Classically simulating the linear folding prescription requires auxiliary qubits to implement the absolute value operation or, much more practically, saving the phase oracle as a pre-computed diagonal operator.
@@
We can also consider trial minimum annealing (TMA), originally proposed in Kapit (2021a). We explore the TMA formulation in depth because we can predict its scaling analytically. In this scheme, a simple classical algorithm is used to find an initial local minimum of $H_{P}$; the quality of the minimum does not particularly matter and for approximation-hard instances we assume it is far above the true ground state energy, in the worst case asymptotically approaching random guessing. Let this classical minimum state be $\left|L\right>$. We will use the linear folding prescription in Eq. 5 for $H_{P}$ itself, and add to it a new, diagonal lowering Hamiltonian $H_{L}$ which has $\left|L\right>$ as its ground state, and assign to it a time-dependent coefficient $C\left(t\right)$. Our total cost function Hamiltonian is
@@
Recall that $H_{P}$ is normalized so that its ground state energy is $-N$. To go further, we need to specify a form for $H_{L}$. For this analysis will choose a new random hypergraph of $N_{C}$ triples which, critically, has no correlation to the hypergraph of $H_{P}$; we choose the same $N_{C}$ as the problem for convenience here but any $O\left(N\right)$ quantity should be fine. We choose the signs of all constraints so that $\left|L\right>$ satisfies all of them. $H_{L}$ is not included in the folding procedure so applied separately in time evolution. We then choose $C\left(t=0\right)$ such that the initial energy of $\left|L\right>$ (defined by $H_{\mathrm{fold}}+C\left(t\right)H_{L}$) is well below $-N$ but remains $O\left(N\right)$. Our algorithm simulates appropriately discretized time evolution in the following sequence:
@@
  • Evolve time for a total time $T$, likely also $O\left(N\right)$, where $C\left(t\right)$ is smoothly ramped down to zero, ensuring that $\left|L\right>$ crosses the hyperspherical shell of ground states of $H_{\mathrm{fold}}=\left|H_{P}+AN\right|/A$. Note that this crossing occurs when the ground state energy of $C\left(t\right)H_{L}$ is $O\left(-N\right)$, and if we assume the initial minimum was uncorrelated with the true ground state $\left|G\right>$, the mean Hamming distance between $\left|L\right>$ and any of the ground states of the folded Hamiltonian is $N/2$ flips. Consequently, $H_{L}$ adds an $O\left(\sqrt{N}\right)$ energy uncertainty to these states that has no meaningful impact on the approximation ratio.
@@
The total gate count of this algorithm is as follows. We have a factor of $O\left(N+N_{C}\;{\rm polylog}\left(N\right)\right)$ per timestep for the layers of transverse field, $H_{\mathrm{fold}}$ and $H_{L}$ terms, which we simplify to $N_{C}\;{\rm polylog\left(N\right)}$. We obtain, in the worst case, a factor of $O\left(N_{C}\right)$ for the number of guesses one needs to make to correctly set the normalization for a chosen $A$. We assume, on empirical grounds, that the total quantum evolution time is $O\left(N\right)$.This choice works well in practice in our simulations, and more intuitively, the very simplest classical optimization routine, steepest descent, requires $O\left(N\right)$ Hamiltonian calls to halt. We do not think it reasonable, ultimately, that a quantum algorithm should perform well with fewer steps per shot. Finally, in the worst case we expect a timestep $dt\sim 1/N$, for graphs where a small number of variables connect to significant fractions of the $N_{C}$ constraints, but $dt$ constant or increasing logarithmically is empirically and intuitively fine in the typical case. Taken together, and we emphasize assuming that the algorithm is capable of returning states with $E<AE_{\mathrm{GS}}$ in constant probability, we estimate a total runtime between $O\left(N_{C}^{2}N^{2}\;{\rm polylog}\left(N\right)\right)$ in the worst case and $O\left(N_{C}N\;{\rm polylog}\left(N\right)\right)$ in more typical cases.
All that said, justifying the assumption of constant success probability is the central task of the paper. We now provide a theoretical analysis of the performance of this formulation of spectral folding on random hypergraph PPSPs, and a more qualitative analysis of the expected performance of other variations.
Figure 3: Illustration of the spectral folding procedure. (Top) Sketch of the rough energy landscape of an approximation-hard CSP, with a single deep minimum whose basin of attraction is an exponentially small fraction of the configuration space. Directly optimizing this cost function through quantum approaches often misses the deep minimum entirely, for reasons explained in the text. (Bottom) Spectrally folded energy landscape in Eq. (5), where the problem Hamiltonian energies are mirrored around an approximation target $E=AE_{\mathrm{GS}}$. This can be implemented in a gate model algorithm with modest overhead, as shown in the text. Doing so promotes the states near the fold to an exponentially large ground state band while eliminating an interference effect that reduces tunneling into them from trivial initial states; for a wide range of problem instances (and likely, low-order problem classes), this works out to an approximation guarantee. Detailed performance predictions, and numerical benchmarking, are shown in the text.
@@ IV Analytical performance predictions
IV.1 Preliminaries
In this section, we will predict from first principles the average macroscopic quantum tunneling rate–and thus, achievable approximation ratio–for the TMA variation of spectrally folded quantum optimization. This version may not be the optimal choice, and there are intuitive reasons to believe that other variations could offer better performance, but being able to make direct analytical predictions enormously strengthens our argument and bolsters the scaling expectations one can infer from our numerical results. To do so, we have developed a somewhat novel resummed extensive order perturbation theory based on previous forward approximation results Bravyi et al. (2011); Pietracaprina et al. (2016); Baldwin et al. (2016, 2017); Scardicchio and Thiery (2017); Baldwin and Laumann (2018); Grattan et al. (2023).
For the random hypergraph problems we study here, the two key factors in determining the macroscopic quantum tunneling rate are the transverse field strength $\kappa_{c}$ where a phase transition occurs between the paramagnet and the quantum spin glass, which we call the dressed problem phase (DPP) in this work, and the energy cost $E\left(x\right)$ for $x$ random flips away from the ground state. We first derive the energy cost, and remarkably, for random flip sequences it turns out to be graph independent. Specifically, for MAX-3-XORSAT, our problem is defined as a hypergraph of $N_{C}$ $p$-body constraints (e.g. $V_{ijk}Z_{i}Z_{j}Z_{k}$) over $N$ variables, where $p=3$ here, and each constraint returns ±$\pm 1$ and flips to the opposite value when any one of the spins flips. Let us say the system is in some classical configuration $s$; the energy is then given by $E\left(s\right)=N_{C}\left(n_{\mathrm{unsat}}-n_{\mathrm{sat}}\right)$, where a sat constraint returns 1 in this notation, and $n$ implies a density.
@@ IV.1 Preliminaries
$\Delta E_{\mathrm{avg}}=+2p\frac{N_{C}}{N}\left(N_{\mathrm{sat}}-N_{\mathrm{unsat}}\right)=-2pE$ (9)
@@
Now imagine we have flipped $y$ spins from our initial configuration. If we flip one more spin at random, once again $\Delta E_{\mathrm{avg}}\left(y\right)=-2pE\left(y\right)\Delta y$. However, we have already flipped $y$ spins, so when we flip one more at random, with probability $1-y/N$ we have flipped a spin back and are computing the energy change associated with reducing $y$ by 1. Consequently
$\left(1-\frac{2y}{N}\right)\frac{\Delta E_{\mathrm{avg}}\left(y\right)}{\Delta y}=-2pE\left(y\right)$ (10)
@@
If we interpret this as a differential equation, it has a straightforward solution: starting from the ground state $G$, for $x$ unique random flips away the average energy is
$E_{\mathrm{avg}}\left(x\right)=E_{\mathrm{GS}}\left(1-\frac{2x}{N}\right)^{p}.$ (11)

### v3

@@ III.1 Key preliminaries – filter functions
A good choice of $A$ is important for spectrally folded quantum optimization to succeed; if (absent warping) $A$ is chosen to be too close to 1, then we risk failing to well-approximate $H_{P}$ due to the interference effects mentioned in section II.3. A choice of $A$ which is too small will return a suboptimal approximation ratio, and if $A$ is too close to zero, cause instabilities from having a poorly defined problem to solve. Fortunately, for random hypergraph MAX-3-XORSAT instances drawn from $\mathcal{P}\left(N_{C},N,\epsilon\right)$ we can predict the threshold $A$ for which we expect a polynomial depth TMA circuit to return states with $E\simeq AE_{\mathrm{GS}}$ from first principles. The ideal value of $A$ depends both on the problem class and on the variation of spectral folding employed; for MAX-3-XORSAT, $A\geq 0.6$ is achievable as derived below in section IV.3.2. This is a significant leap over the best known classical approximation algorithm for this problem Barak et al. (2015), which offers a weaker guarantee with much more restricted viability, in the worst case, equivalent to $A\to 0$ in our notation. It is also a significant leap over recent quantum approaches to this problem Anshu et al. (2021); Marwaha and Hadfield (2022). We note that minimizing $\left(H-E\right)^{2}$ is not itself a novel idea and has been used in classical and quantum algorithms for finding states close to specific energies in chemical and many-body systems Wang and Zunger (1994a, b); McClean et al. (2016); Santagati et al. (2018); Zhang et al. (2021); Tilly et al. (2022); Tazi and Thom (2023). To our knowledge, however, the use of spectral folding for approximate optimization of CSPs is novel, both in concept and in the analysis we present below to choose $A$ and understand at a deeper level why it presents significant advantages over optimizing the problem $H_{P}$ directly. We similarly are not aware of any previous work employing non-quadratic folding, or where a warping transformation as in Eq. 9 is applied to optimize a classical CSP.
@@
From hereon, we let $H_{\mathrm{fold}}$ be the spectrally folded problem Hamiltonian. Having defined it, there are a number of ways we can attempt to find its ground states. We present three – first, for completeness, the direct conventional approach, inspired by AQC, and then two new protocols.
III.2 Direct method
The simplest choice, and one that performs well empirically, is AQC-inspired state preparation, where we interpolate in Trotterized evolution between the transverse field driver and the folded problem over a time $t_{F}$:
@@ III.2 Direct method
To simulate this prescription classically, it is easiest to simply construct the phase oracle as a diagonal operator applied to the $N$ qubit state at each timestep. As we shall show later, with $x=2$, $A=0.75$, and $t_{F}$ growing linearly with $N$, this algorithm performs well in practice with numerical simulation, though we cannot predict its performance analytically. To achieve a better approximation of the ground state–and, crucially, a formulation where performance can be predicted from first principles–we need to apply a warping filter as well, or consider tunneling between a spin glass minimum and the folded ground state band.
III.3 Multi-stage filtered optimization
We now present a more complex, and more powerful, formulation that we call multi-stage filtered optimzation (MSFO). We will first state the steps of the algorithm, then in section IV.2 provide a detailed explanation and analysis of its stages, and expected performance. As mentioned earlier, we assume that we know the fraction $\epsilon$ of clauses which are unsatisfied in the ground state, so we can properly set the normalization of $H_{P}$ so that the ground state energy is $-N$ and random states have energy zero. In the worst case, there is an overhead of $O\left(N_{C}\right)$ for guessing this value; the analysis below proceeds for the trials where it was guessed correctly. MSFO follows the following sequence, per-shot:
  • First, we initialize the uniform superposition $\left|S\right>$, the ground state of $H_{D}=-\kappa\sum_{j}X_{j}$; here $\kappa$ is the transverse field strength. We use $\kappa=1$ for simulations later in this work but other $O\left(1\right)$ choices can of course be used. For the first stage, we optimize the following cost function, comprised of two composed spectral filters defined in Eqs. (8,9) applied successively to $H_{P}$ (see FIG. 4 for an illustration), with parameters $x\geq 2$ and $w<1$:
    $H_{\mathrm{cost}}^{x,w}$ $=$ $\sum_{n=0}^{2^{N}-1}F_{w}\left(E_{f}\left(n\right),N,w\right)\left|n\right>\left<n\right|,$ (12)
    $E_{f}\left(n\right)$$\equiv$ $F_{f}\left(E\left(n\right),N,1,x\right).$
  • We then choose a time $T_{1}$ and timestep $dt$ for the first stage, as well as a function $s\left(t\right)$ which ramps from 0 to 1 over the time $T_{1}$. In this work, we choose the simplest possible choice $s\left(t\right)=t/T_{1}$, but undoubtedly benefits can be gained by further tuning. In this stage, we evolve time in Trotterized evolution from $t=0$ to $t=T_{1}$ by iterating
    $\left|\psi\left(t+dt\right)\right>=e^{-2\pi idt\kappa H_{D}}e^{-2\pi idts\left(t\right)H_{\mathrm{cost}}^{x,w}}\left|\psi\left(t\right)\right>.$ (13)
  • We have now concluded stage 1. For stage 2, we pick a new duration $T_{2}$, and two new functions $s_{x}\left(t\right)$ and $s_{w}\left(t\right)$. These functions are used to ramp off the nonlinear folding and warping transformations, so that at the end of stage 2, the classical cost function is physically equivalent to the un-filtered $H_{P}$. Both of these functions are defined such that $s_{x}\left(0\right)=1$, $s_{x}\left(T_{2}\right)=1/x$ and $s_{w}\left(0\right)=1$, $s_{w}\left(T_{2}\right)=1/w$. We evolve time from $t=0$ to $t=T_{2}$ by iterating:
    $\left|\psi\left(t+dt\right)\right>=e^{-2\pi idt\kappa H_{D}}e^{-2\pi idtH_{\mathrm{cost}}^{xs_{x}\left(t\right),ws_{w}\left(t\right)}}\left|\psi\left(t\right)\right>.$ (14)
  • In practice, linear interpolation is completely sufficient for the second stage. The nonlinear folding and warping transformations have now been smoothly turned off. For our final stage, we pick a third evolution time $T_{3}$, and evolve time as in the previous stages, this time ramping down the transverse field strength $\kappa$ to zero. We then measure the system in the computational basis, and this shot is concluded.
The simple linear schedules mentioned above are not optimal, but they do not need to be. In the next section we show that, for $H_{P}$ as defined in our core conjecture, $T_{1}$ increasing subquadratically with $N$, and provided that we can ignore transverse field chaos, e.g. corrections pushing local minima of $H_{\mathrm{cost}}^{x,w}$ below $G$ and consequently steering the system toward those states first, evolution in the first stage will find the quantum ground state of $H_{D}+H_{\mathrm{cost}}^{x,w}$ with $O\left(1\right)$ probability. The runtime of this stage is set by the scaling of a single transition that occurs when $E_{0,D}\simeq-\kappa N$ crosses $E_{G,D}$, the energy of the dressed ground state $\left|G_{D}\right>$ of $H_{\mathrm{cost}}^{x,w}$, and the minimum gap scales proportionally to $S_{D}|G_{D}\right>$, where $\left|S_{D}\right>$ is ground state of $H_{D}$ as dressed by corrections from $H_{\mathrm{cost}}^{x,w}$. As discussed in section IV.2, the first stage forms the bottleneck of the algorithm and subsequent stages are much easier, both to understand conceptually and in their minimum gate count. The asymptotic gate count of this algorithm depends on the choice of $x$ and $w$ and is considered in section IV.2.
III.4 Spectrally folded trial minimum annealing
@@
For our third and final algorithm, we consider trial minimum annealing (TMA), originally proposed in Kapit (2021a). We explore the TMA formulation in depth because we can predict its scaling analytically. In this scheme, a simple classical algorithm is used to find an initial local minimum of $H_{P}$; the quality of the minimum does not matter and for approximation-hard instances we assume it is far above the true ground state energy, in the worst case asymptotically approaching random guessing. Let this classical minimum state be $\left|L\right>$. We will use the linear folding prescription ($x=1$, no warping transformation applied) in Eq. 8 for $H_{P}$ itself, and add to it a new, diagonal lowering Hamiltonian $H_{L}$ which has $\left|L\right>$ as its ground state, and assign to it a time-dependent coefficient $C\left(t\right)$. Finally, we choose an approximation target $A\leq 1$ that specifies the energy $-AN$ (where $E_{GS}=-N$ is negative in our conventions) that our algorithm aims to find states near. Our total cost function Hamiltonian is
@@
This folding form is a simplification of the generalized nonlinear folding function Eq. 8, with $x=1$; the symmetrization there is unnecessary in practice for the values of $A\geq 0.6$ considered in this work.
To go further, we need to specify a form for $H_{L}$. For this analysis will choose a new random hypergraph of $N_{C}$ triples which, critically, has no correlation to the hypergraph of $H_{P}$; we choose the same $N_{C}$ as the problem for convenience here but any $O\left(N\right)$ quantity should be fine. We choose the signs of the new constraints so that $\left|L\right>$ satisfies all of them. $H_{L}$ is not included in the folding procedure so applied separately in time evolution. We then choose $C\left(t=0\right)$ such that the initial energy of $\left|L\right>$ (defined by $H_{\mathrm{fold}}+C\left(t\right)H_{L}$) is well below $-N$ but remains $O\left(N\right)$. Our algorithm simulates appropriately discretized time evolution in the following sequence:
@@
  • Evolve time for a total time $T$, likely also $O\left(N\right)$, where $C\left(t\right)$ is smoothly ramped down to zero, ensuring that $\left|L\right>$ crosses the hyperspherical shell of ground states of $H_{\mathrm{fold}}=\left|H_{P}+AN\right|/A$. Note that this crossing occurs when the ground state energy of $C\left(t\right)H_{L}$ is $O\left(-N\right)$, and if we assume the initial minimum was uncorrelated with the true ground state $\left|G\right>$, the mean Hamming distance between $\left|L\right>$ and any of the ground states of the folded Hamiltonian is $N/2$ flips. Consequently, $H_{L}$ adds an $O\left(\sqrt{N}\right)$ energy uncertainty to these states that has no meaningful impact on the approximation ratio. The algorithm succeeds if, at this stage, the many-body state transitions into any state near the fold energy through a collective tunneling process, and we predict the dynamics and scaling of this stage in the next section.
@@
III.5 Resource estimates for spectrally folded trial minimum annealing
The total gate count of this algorithm is as follows. We have a factor of $O\left(N+N_{C}\;{\rm polylog}\left(N\right)\right)$ per timestep for the layers of transverse field, $H_{\mathrm{fold}}$ and $H_{L}$ terms, which we simplify to $N_{C}\;{\rm polylog\left(N\right)}$. We obtain, in the worst case, a factor of $O\left(N_{C}\right)$ for the number of guesses one needs to make to correctly set the normalization for a chosen $A$. We assume, on empirical grounds, that the total quantum evolution time is $O\left(N\right)$, for the reasons discussed in section II.5. Finally, in the worst case we expect $dt$ may need to decrease polynomially at large $N$, but $dt$ constant or increasing logarithmically is empirically and intuitively fine in the typical case (see also Granet and Dreyer (2024)). Taken together, and we emphasize assuming that the algorithm is capable of returning states with $E<AE_{\mathrm{GS}}$ in constant probability, we estimate a total runtime between $O\left(N_{C}^{2}N^{2}\;{\rm polylog}\left(N\right)\right)$ in the worst case and $O\left(N_{C}N\;{\rm polylog}\left(N\right)\right)$ in more typical cases. Assuming constant per-shot success probability, the asymptotic runtimes of MSFO and TMA are thus similar. Justifying that assumption for both algorithms is in some sense the core task of this paper, which we now begin.
Figure 3: Illustration of the spectral folding procedure. (Top) Sketch of the rough energy landscape of an approximation-hard CSP, with a single deep minimum whose basin of attraction is an exponentially small fraction of the configuration space. Directly optimizing this cost function through quantum approaches often misses the deep minimum entirely, due to mechanisms proposed in section II and confirmed numerically in section V.3. (Bottom) Spectrally folded energy landscape in Eq. (8), where the problem Hamiltonian energies are mirrored around an approximation target $E=AE_{\mathrm{GS}}$. This can be implemented in a gate model algorithm with modest overhead, as shown in the text. Doing so promotes the states near the fold to an exponentially large ground state band while eliminating an interference effect that reduces tunneling into them from trivial initial states; for a wide range of problem instances (and likely, low-order problem classes), this works out to an approximation guarantee. Detailed performance predictions, and numerical benchmarking, are shown in the text.
@@ IV Analytical performance predictions
We are now ready to present the main results of this paper: analytical performance predictions for MSFO and spectrally folded TMA. For all the reasons that we specified in the previous sections, we consider approximate optimization to be a much more promising objective for achieving exponential speedups with quantum algorithms than attempts to directly find the exact ground state of hard CSPs. But this too would seem to be an enormously difficult task, given the complexity of performing a rigorous calculation for any individual disordered problem hypergraph. We make this challenge tractable through the following methods. We first narrow our focus to random instances drawn from the space of those which are formally approximation-hard, which allows us to take advantage of certain statistical properties of these random hypergraphs. Second, we present two algorithmic approaches which, by design, are constructed to minimize the influence of specific details of any given instance on final results.
The two approaches are very different. MSFO uses a sequence of composed filtering transformations to overcome the first order transition that makes more traditional attempts to find the ground state with quantum time evolution take exponential time, at the cost of potentially worsening TFC. For this algorithm we can predict the time to solution scaling through analogy to mean field models, a claim we subsequently verify in numerical simulation. In contrast, TMA applies a folding transformation to promote an exponentially large band of excited states to ground states of the transformed cost function, and then finds them through collective tunneling, at the cost of being unable to find any lower energy states directly. To predict the scaling and parameter choices of this algorithm, we have developed a somewhat novel resummed extensive order perturbation theory based on previous forward approximation results Bravyi et al. (2011); Pietracaprina et al. (2016); Baldwin et al. (2016, 2017); Scardicchio and Thiery (2017); Baldwin and Laumann (2018); Grattan et al. (2023).
IV.1 Cost per flip curve
Underpinning both the formulation of our novel algorithms, and our analysis of their performance, is an asymptotically exact, hypergraph-independent Hastings (2018) prediction of the average energy $E_{\mathrm{avg}}\left(x\right)$ for $x$ random classical bit flips away from the ground state. Specifically, for MAX-3-XORSAT, our problem is defined as a hypergraph of $N_{C}$ $p$-body constraints (e.g. $V_{ijk}Z_{i}Z_{j}Z_{k}$) over $N$ variables, where $p=3$ here, and each constraint returns ±$\pm 1$ and flips to the opposite value when any one of the spins flips. Let us say the system is in some classical configuration $s$; the energy is then given by $E\left(s\right)=N_{C}\left(n_{\mathrm{unsat}}-n_{\mathrm{sat}}\right)$, where a sat constraint returns 1 in this notation, and $n$ implies a density.
@@ IV.1 Cost per flip curve
$\Delta E_{\mathrm{avg}}=+2p\frac{N_{C}}{N}\left(N_{\mathrm{sat}}-N_{\mathrm{unsat}}\right)=-2pE.$ (16)
@@
This is an exact statement, though again it applies only to averages. Now imagine we have flipped $x$ spins from our initial configuration. If we flip one more spin at random, once again $\Delta E_{\mathrm{avg}}\left(x\right)=-2pE\left(x\right)\Delta x$. However, we have already flipped $x$ spins, so when we flip one more at random, with probability $x/N$ we have flipped a spin back and are computing the energy change associated with reducing $x$ by 1. Consequently
$\left(1-\frac{2x}{N}\right)\frac{\Delta E_{\mathrm{avg}}\left(x\right)}{\Delta x}=-2pE_{\mathrm{avg}}\left(x\right).$ (17)
@@
If we combine this with $E_{\mathrm{avg}}\left(0\right)=E\left(G\right)$ and $E_{\mathrm{avg}}\left(N/2\right)=0$ (as $N/2$ flips reaches fully random and uncorrelated configurations), we can interpret this as a differential equation discretized over $N/2$ steps. This differential equation has a straightforward solution: starting from the ground state $G$, for $x$ unique random flips away the average energy is
$E_{\mathrm{avg}}\left(x\right)=E\left(G\right)\left(1-\frac{2x}{N}\right)^{p}.$ (18)